40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 3 12:27:01 2011, Nicole, Summary, SUS, Weekly Summary calibrationplot.jpgcalibrationplot.jpgcalibrationplot.jpgnewholestobe.JPG
    Reply  Wed Aug 3 12:37:57 2011, Koji, Summary, SUS, Weekly Summary 
Message ID: 5107     Entry time: Wed Aug 3 12:27:01 2011     Reply to this: 5108
Author: Nicole 
Type: Summary 
Category: SUS 
Subject: Weekly Summary 

This week I have determined the linear region for my photosensor. I have determined the linear region to be -14.32 V/cm in the region 0.4cm 0.75 cm.

In order to obtain this voltage plot, I used a 287K resistor to set the max voltage output for the photodiodes. This calibration was obtained using a small rectangular standing mirror (not the TT testing mirror that Steve has ordered for me).

calibrationplot.jpg

I have also been working on the second half of the photosensor circuit (to power the LED and read out voltages for the second photosensor head). I have assembled the constant-current section of the circuit and need to do the voltage-output section of the circuit. I also need to finish assembling the second photosensor head and cables.

 

I submitted my Second Progress Report on Tuesday.

 

I have attached the mirror to the TT suspension. We are using 0.006 diameter tungsten wire to suspend the mirror. I am currently working on balancing the mirror.

 

This morning, I realized that the current set-up of the horizontal shaker does not allow for the TT to be securely mounted. I was going to change the drill holes in the horizontal slider base (1 inch pitch). Jamie has suggested that it is better to make a pair of holes in the base larger. The circled holes are the ones that will be expanded to a 0.26" diameter so that I can mount the mirror securely to the horizontal slider base. There is a concern that a bit of the TT suspension base will hang over the edge of the horizontal sliding plate. We are not sure if this will cause problems with shaking the mirror evenly. Suggestions/advice are appreciated.

newholestobe.JPG

Attachment 1: calibrationplot.jpg  34 kB  | Hide | Hide all
calibrationplot.jpg
Attachment 2: calibrationplot.jpg  34 kB  | Hide | Hide all
calibrationplot.jpg
ELOG V3.1.3-